# Uncertainty Propagation, and Parameter Calibration

### v1

@@ Abstract
The goal of this paper is to demonstrate and address challenges related to all aspects of performing a complete uncertainty quantification (UQ) analysis of a complicated physics-based simulation like a 2D slab burner direct numerical simulation (DNS). The UQ framework includes the development of data-driven surrogate models, propagation of parametric uncertainties to the fuel regression rate–the primary quantity of interest–and Bayesian calibration of critical parameters influencing the regression rate using experimental data. Specifically, the parameters calibrated include the latent heat of sublimation and a chemical reaction temperature exponent. Two surrogate models, a Gaussian Process (GP) and a Hierarchical Multiscale Surrogate (HMS) were constructed using an ensemble of 64 simulations generated via Latin Hypercube sampling. Both models exhibited comparable performance during cross-validation. However, the HMS was more stable due to its ability to handle multiscale effects, in contrast with the GP which was very sensitive to kernel choice. Analysis revealed that the surrogates do not accurately predict all spatial locations of the slab burner as-is. Subsequent Bayesian calibration of the physical parameters against experimental observations resulted in regression rate predictions that closer align with experimental observation in specific regions. This study highlights the importance of surrogate model selection and parameter calibration in quantifying uncertainty in predictions of fuel regression rates in complex combustion systems.
@@ 1 Introduction
There are many methods for developing surrogate models. Polynomial chaos expansions (PCE) approximate the input to QoI function $\mathcal{M}:\bm{x}\in\mathcal{D}_{\bm{X}}\subset\mathbb{R}^{M}\mapsto\bm{y}=\mathcal{M}(\bm{x})$ with polynomial approximations [27]. Canonical low-rank tensor approximations (LRA) approximate the input to QoI function with products of univariate functions that can also be orthogonal polynomials similar to PCE [28]. Feedforward neural networks map inputs to QoIs via learning the weights and biases between layers of neurons [29, 30]. Kriging or Gaussian Processes (GPs) assume the likelihood function of the QoI to be a multivariate normal distribution modeled as the addition of a mean regression term and a Gaussian stationary random process [31]. Combinations of these methods are also options for surrogate models, e.g., Deep GPs [32], where the data between layers of a neural network are modeled as a multivariate GP. Traditional surrogate methods such as the ones we outline above often perform poorly when used in multiscale problems, because these problems include characteristics that are computationally challenging for these basic methods. Firstly, multiscale problems have a very large input space and that requires specific techniques to address the curse of dimensionality before proceeding with surrogate models [33, 34, 35]. Secondly, methods such as GPs are not scalable because they include covariance matrix inversions, which are computationally expensive. Thirdly, the assumptions of smoothness and stationarity imposed by Gaussian assumptions or polynomials does not necessarily hold for multiscale problems. Lastly, the ability of simpler methods to capture non-linear trends from high-dimensional data is limited [36].
@@
The main goal of this paper is to systematically address challenges related to all aspects of performing a complete UQ analysis of a complicated physics-based simulation like a 2D slab burner DNS. We provide insights related to the development of an acceptable surrogate model, the propagation of uncertainty of the DNS inputs to the QoI, and the calibration of DNS parameters based on observations of experimental data. The DNS used in this work is ABLATE (Ablative Boundary Layers At The Exascale), an in-house DNS solver developed for simulating a multiphase 2D slab burner setup to better understand combustion in hybrid rocket motors, which includes ablation, radiation, and soot modeling[37, 21, 15]. ABLATE uses the PETSc libraries from Argonne National Laboratory [38, 39] for data management, and parallel large scale computation. For the work described in this paper, ABLATE simulated the 2D reacting flow between pure oxygen $O_{2}$ (the oxidizer) and Methyl methacrylate MMA (the fuel) with chemistry modeled by a reduced mechanism. To accurately capture the multiscale behavior and overcome the computationally prohibitive UQ analysis of the DNS simulations, we build both GPs and a novel computationally efficient hierarchical multiscale surrogate (HMS) [40]. HMS is based on a forward-backward greedy approach to construct a set of data driven basis functions with a multiscale structure. Forward-backward here means the iterative selection of basis functions at different scales. This approach generates hierarchical basis functions belonging to Reproducing Kernel Hilbert Spaces (RKHS), and the hierarchical models are able to accurately represent irregularly structured data at different resolutions (scales), rapidly minimizing the error in fitting by using kernels at successively finer scales. We conduct a forward UQ analysis using the developed surrogate and uncertainty in a subset of inputs identified using sensitivity analysis and propagating the uncertainty of the DNS inputs to the slab burner output QoI. The inputs to the DNS are the velocity factor for the far-field inlet oxidizer flux assuming a parabolic profile ($V_{in}$), the thermochemical properties of the chemistry mechanism ($\theta_{c}$: activation energies $E_{*}$, temperature exponents $b_{*}$, and pre-exponential factors $A_{*}$), the fixed geometry of the slab burner, and the fuel parameters ($\theta_{f}$). Combustion DNS solvers produce a variety of outputs that may be of interest depending on the goals. Here, we are most interested in QoIs along the boundary of the slab, namely the regression rate ($\dot{r}$). The regression rate has a direct impact on the geometrical design of the rocket motor and its performance ([41], [42]). Lastly, we leverage the constructed surrogate model and experimental regression rate measurements to calibrate the fuel ($\theta_{f}$) and chemistry ($\theta_{c}$) parameters of the DNS. In particular, we use the Bayesian method to infer the probability distribution of the latent heat of sublimation ($l_{v}$) and temperature exponent for a particular important reaction ($b_{249}$).
@@ 2 ABLATE: Coupled Flow/Combustion DNS
In this section, we describe our coupled flow/combustion solver ABLATE (https://ablate.dev/), its inputs and outputs, and how we used it to generate the dataset to train the surrogates in Section 3. The solver simulates a 2D slab burner setup for the combustion of MMA as the fuel and pure oxygen $O_{2}$ as the oxidizer. Figure  1 shows the two-dimensional slab burner domain. An inlet of pure oxygen flows from the left side over the fuel slab where it reacts with the fuel vapors and flows through the outlet of the slab burner on the right side. There are four different boundary conditions used in the simulation depending if the boundary is a (1) wall, (2) inlet, (3) outlet, or (4) the fuel. The gas phase system for the slab burner environment is described by the reacting compressible Navier-Stokes equations, coupled with radiation heat transfer:
$\frac{\partial\rho}{\partial t}+\frac{(\partial\rho u_{i})}{\partial x_{i}}=0$
$\frac{\partial\rho u_{j}}{\partial t}+\frac{\partial(\rho u_{i}u_{j})}{\partial x_{i}}=-\frac{\partial p}{\partial x_{j}}+\frac{\partial\tau_{ij}}{\partial x_{i}}$
$\frac{\partial(\rho Y_{k})}{\partial t}+\frac{\partial(\rho u_{i}Y_{k})}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left(\rho D_{k}\frac{\partial Y_{k}}{\partial x_{i}}\right)+\dot{m}^{\prime\prime\prime}_{k}$ (1)
$\frac{\partial(\rho e_{t})}{\partial t}+\frac{\partial(\rho u_{i}H_{t})}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left(u_{i}\mu\frac{\partial u_{i}}{\partial x_{i}}\right)+\frac{\partial}{\partial x_{i}}\left[\left(k\frac{\partial T}{\partial x_{i}}+\sum^{N}\rho D_{k}\frac{\partial Y_{k}}{\partial x_{i}}h_{k}\right)\right]$
@@
$\hskip 91.04872pt-\sum\dot{m}^{\prime\prime\prime}_{k}h_{f,k}^{\circ}-\frac{\partial}{\partial x_{i}}(\rho Y_{C(s)}V_{T,i}h_{C(s)})-\kappa_{P}(\sigma T^{4}-G)$
where, $k$ is the thermal conductivity, $\dot{m}^{\prime\prime\prime}_{k}$ is the mass consumption or production rate of the $k^{th}$ species from chemical reactions, $e_{t}=e+u_{i}^{2}/2$ is the total sensible energy, $H_{t}=e_{t}+p/\rho$ is the total enthalpy and $h_{f,k}^{\circ}$ is the heat of formation of the $k^{th}$ species. $V_{t,x}$ is the thermophoretic velocity associated with soot thermophoresis. A SutherLand transport model is used along with a unity Lewis number assumption. The species diffusivities are assumed to be the same for all species. The chemical kinetic mechanism for MMA oxidation, developed by Bolshova et. al. [43], is utilized consisting of $67$ species and $263$ reactions. A six step semi-empirical soot model based on an acetylene precursor is adapted that includes rates for nucleation, agglomeration, surface growth and oxidation [44].
@@
The wall boundary conditions are simply treated as isothermal no-slip walls held at the ambient temperature of 300 K. The inlet and outlet boundary conditions are described using locally one-dimensional inviscid characteristics (LODI) [48]. The velocity profile prescripted at the inlet is described using a $1/7^{th}$ power law for fully developed turbulent profile for pipe flow, $u_{in}(r)=V_{in}(1-2r/D_{pipe})^{1/7}$, where $V_{in}$ is a velocity factor defined by the oxidizer mass flux to obtain the correct mass flow rate, $A=60~{}G/49\rho_{ox}$.
@@
where $\dot{q}^{\prime\prime}_{c,g}(=k\vec{\nabla}T\cdot\hat{n}_{s})$ is the conductive heat flux from the gas phase normal to the fuel surface, $\dot{q}^{\prime\prime}_{c,f}$ is the conductive heat flux used in heating the fuel, and $\dot{q}^{\prime\prime}_{r,Net}(=\alpha_{s}\dot{q}^{\prime\prime}_{r,abs}-\epsilon_{s}\sigma T_{f}^{4}$) is the net radiation heat transfer at the fuel surface. $\dot{q}^{\prime\prime}_{r,abs}$ is the irradiant radiation heat flux onto the fuel surface, and $T_{f}$ is the fuel surface temperature. $\epsilon_{s}$ and $\alpha_{s}$ are the fuel surface emissivity and absorptivity. $\dot{m}^{\prime\prime}_{f}=\rho_{f}\dot{r}$ is the local fuel mass flux and is related to the local regression rate by the solid fuel density $\rho_{f}$. The current simulations assume the fuel is isothermal and already at the vaporization temperature of PMMA ($T_{f}=653~{}K$) such that the is no heat loss due to heating the fuel ($\dot{q}_{c,f}=0$). This leads to the following mass flux and regression rate relations depending only on the instantaneous gas phase solution,
@@
In this isothermal fuel boundary model, the fuel temperature is held at $T_{f}=653K$. The density, MMA species mass, and normal momentum boundary conditions have additional mass and momentum fluxes due to $\dot{m}^{\prime\prime}_{f}$.
@@
Where $\mathcal{S}_{B}$ includes the paired coordinates $x,y$ for the grid points on the boundary, $N$ is the number of simulation time steps, $t_{i}$ is the elapsed simulation time, and $\dot{r}(x,y,t)$ is the regression rate as computed during the simulation from Eqn. 3. Each simulation of the DNS resolves, on average, about 25-75ms of the reacting flow (1 to 3 flow through times) and takes 24 hours to complete.
3 Surrogate Models and Uncertainty Quantification Methods

### v2

# UQ of 2D Slab Burner DNS: Surrogates, Uncertainty Propagation, and Parameter Calibration

@@ Abstract
The goal of this paper is to demonstrate and address challenges related to all aspects of performing a complete uncertainty quantification analysis of a complicated physics-based simulation like a 2D slab burner direct numerical simulation (DNS). The UQ framework includes the development of data-driven surrogate models, propagation of parametric uncertainties to the fuel regression rate–the primary quantity of interest–and Bayesian calibration of the latent heat of sublimation and a chemical reaction temperature exponent using experimental data. Two surrogate models, a Gaussian Process (GP) and a Hierarchical Multiscale Surrogate (HMS) were constructed using an ensemble of 64 simulations generated via Latin Hypercube sampling. HMS is superior for prediction demonstrated by cross-validation and able to achieve an error $<15\%$ when predicting multiscale boundary quantities just from a few far field inputs. Subsequent Bayesian calibration of chemical kinetics and fuel response parameters against experimental observations showed that the default values used in the DNS should be higher to better match measurements. This study highlights the importance of surrogate model selection and parameter calibration in quantifying uncertainty in predictions of fuel regression rates in complex combustion systems.
@@ 1 Introduction
There are many methods for developing surrogate models. Polynomial chaos expansions (PCE) approximate the input to QoI function $\mathcal{M}:\bm{x}\in\mathcal{D}_{\bm{X}}\subset\mathbb{R}^{M}\mapsto\bm{y}=\mathcal{M}(\bm{x})$ with polynomial approximations [27]. Canonical low-rank tensor approximations (LRA) approximate the input to QoI function with products of univariate functions that can also be orthogonal polynomials similar to PCE [28]. Feedforward neural networks map inputs to QoIs via learning the weights and biases between layers of neurons [29, 30]. Kriging or Gaussian Processes (GPs) assume the likelihood function of the QoI to be a multivariate normal distribution modeled as the addition of a mean regression term and a Gaussian stationary random process [31]. Combinations of these methods are also options for surrogate models, e.g., Deep GPs [32], where the data between layers of a neural network are modeled as a multivariate GP. Traditional surrogate methods such as the ones we outline above often perform poorly when used in multiscale problems, because these problems include characteristics that are computationally challenging for these basic methods. Firstly, multiscale problems have a very large input space and that requires specific techniques to address the curse of dimensionality before proceeding with surrogate models [33, 34]. Secondly, methods such as GPs are not scalable because they include covariance matrix inversions, which are computationally expensive. Thirdly, the assumptions of smoothness and stationarity imposed by Gaussian assumptions or polynomials does not necessarily hold for multiscale problems. Lastly, the ability of simpler methods to capture non-linear trends from high-dimensional data is limited [35].
@@
The main goal of this paper is to systematically address challenges related to all aspects of performing a complete UQ analysis of a complicated physics-based simulation like a 2D slab burner DNS. We provide insights related to the development of an acceptable surrogate model, the propagation of uncertainty of the DNS inputs to the QoI, and the calibration of DNS parameters based on observations of experimental data. The DNS used in this work is ABLATE (Ablative Boundary Layers At The Exascale), an in-house DNS solver developed for simulating a multiphase 2D slab burner setup to better understand combustion in hybrid rocket motors, which includes ablation, radiation, and soot modeling[36, 21, 15]. ABLATE uses the PETSc libraries from Argonne National Laboratory [37] for data management, and parallel large scale computation. For the work described in this paper, ABLATE simulated the 2D reacting flow between pure oxygen $O_{2}$ (the oxidizer) and Methyl methacrylate MMA (the fuel) with chemistry modeled by a reduced mechanism. To accurately capture the multiscale behavior and overcome the computationally prohibitive UQ analysis of the DNS simulations, we build both GPs and a novel computationally efficient hierarchical multiscale surrogate (HMS) [38]. HMS is based on a forward-backward greedy approach to construct a set of data driven basis functions with a multiscale structure. Forward-backward here means the iterative selection of basis functions at different scales. This approach generates hierarchical basis functions belonging to Reproducing Kernel Hilbert Spaces (RKHS), and the hierarchical models are able to accurately represent irregularly structured data at different resolutions (scales), rapidly minimizing the error in fitting by using kernels at successively finer scales. We conduct a forward UQ analysis using the developed surrogate and uncertainty in a subset of inputs identified using sensitivity analysis and propagating the uncertainty of the DNS inputs to the slab burner output QoI. The inputs to the DNS are the velocity factor for the far-field inlet oxidizer flux assuming a parabolic profile ($V_{in}$), the thermochemical properties of the chemistry mechanism ($\theta_{c}$: activation energies $E_{*}$, temperature exponents $b_{*}$, and pre-exponential factors $A_{*}$), the fixed geometry of the slab burner, and the fuel parameters ($\theta_{f}$). Combustion DNS solvers produce a variety of outputs that may be of interest depending on the goals. Here, we are most interested in QoIs along the boundary of the slab, namely the regression rate ($\dot{r}$). The regression rate has a direct impact on the geometrical design of the rocket motor and its performance ([39], [40]). Lastly, we leverage the constructed surrogate model and experimental regression rate measurements to calibrate the fuel ($\theta_{f}$) and chemistry ($\theta_{c}$) parameters of the DNS. In particular, we use the Bayesian method to infer the probability distribution of the latent heat of sublimation ($l_{v}$) and activation energy for a particular important reaction ($E_{1}$).
@@ 2 ABLATE: Coupled Flow/Combustion DNS
In this section, we describe our coupled flow/combustion solver ABLATE (https://ablate.dev/), its inputs and outputs, and how we used it to generate the dataset to train the surrogates in Section 3. The solver simulates a 2D slab burner setup for the combustion of MMA as the fuel and pure oxygen $O_{2}$ as the oxidizer.
@@
Figure  1 shows the two-dimensional slab burner domain. An inlet of pure oxygen flows from the left side over the fuel slab where it reacts with the fuel vapors and flows through the outlet of the slab burner on the right side. There are four different boundary conditions used in the simulation depending if the boundary is a (1) wall, (2) inlet, (3) outlet, or (4) the fuel. The gas phase system for the slab burner environment is described by the reacting compressible Navier-Stokes equations coupled with radiation heat transfer:
$\frac{\partial\rho}{\partial t}+\vec{\nabla}\cdot(\rho\vec{u})=0$
$\frac{\partial\rho\vec{u}}{\partial t}+\vec{\nabla}\cdot(\rho\vec{u}\vec{u})=-\vec{\nabla}P+\vec{\nabla}\cdot\bm{\tau}^{T}$
$\frac{\partial(\rho Y_{k})}{\partial t}+\vec{\nabla}\cdot(\rho\vec{u}Y_{k})=\vec{\nabla}\cdot\left(\rho D\vec{\nabla}Y_{k}\right)+\dot{m}^{\prime\prime\prime}_{k}$ (1)
$\frac{\partial(\rho E)}{\partial t}+\vec{\nabla}\cdot(\rho\vec{u}H_{t})=\vec{\nabla}\cdot\left(\underline{\underline{\tau}}\cdot\vec{u}\right)+\vec{\nabla}\cdot\left(k\vec{\nabla}T+\sum_{k=1}^{N_{sp}}\rho Dh_{k}\vec{\nabla}Y_{k}\right)$
@@
$\hskip 91.04872pt-\sum_{k=1}^{N_{sp}}\dot{m}^{\prime\prime\prime}_{k}h_{f,k}^{\circ}-\vec{\nabla}\cdot\left(\rho Y_{C(s)}\vec{V}_{T}h_{C(s)}\right)-\kappa_{P}(\sigma T^{4}-G)$
where, $k$ is the thermal conductivity, and $E=e+\vec{u}\cdot{\vec{u}}/2$ is the total sensible energy. $Y_{k}$ are the species mass fractions. $\underline{\underline{\tau}}~{}=\mu(\vec{\nabla}\vec{u}+(\vec{\nabla}\vec{u})^{T}-2/3~{}\underline{\underline{I}}(\vec{\nabla}\cdot\vec{u}))$ is the viscous stress tensor, $H_{t}=E+p/\rho$ is the total sensible enthalpy and $h_{f,k}^{\circ}$ is the heat of formation of the $k^{th}$ species calculated using NASA 7 coefficients [41]. An ideal gas equation of state is used to define the pressure, $P=\rho R_{u}T/MW$, with universal gas constant, $R_{u}$ (=8314.459 $J/kmol-K$) and mixture molecular weight, $MW=(\sum_{k=1}^{N_{sp}}Y_{k}/MW_{k})^{-1}$, where $MW_{k}$ is the molecular weight of the $k^{th}$ species.
$\dot{m}^{\prime\prime\prime}_{k}$ is the mass consumption or production rate of the $k^{th}$ species from chemical reactions defined in compact notation as,
$\dot{m}^{\prime\prime\prime}_{k}=MW_{k}\sum_{r=1}^{N_{rxn}}(\nu_{k,r}^{\prime\prime}-\nu_{k,r}^{\prime})\dot{q}_{r},$ (2)
where $\nu_{k,r}^{\prime\prime}$ and $\nu_{k,r}^{\prime}$ are the stoichiometric coefficients on the products and reactants side of the $k^{th}$ species for reaction $r$. $\dot{q}_{r}$ is the reaction rate of progress variable for reaction $r$ defined as,
$\dot{q}_{r}=k_{F,r}\Pi_{k=1}^{N_{sp}}\left[\frac{\rho Y_{k}}{MW_{k}}\right]^{\nu_{k,r}^{\prime}}-k_{R,r}\Pi_{k=1}^{N_{sp}}\left[\frac{\rho Y_{k}}{MW_{k}}\right]^{\nu_{k,r}^{\prime\prime}},$ (3)
where $k_{F,r}$ and $k_{R,r}$ are the forward and reverse rate constants of reaction $r$ respectively. In this study, we used a detailed chemical kinetic mechanism for MMA oxidation developed by Bolshava et. al. [42], which consists of $N_{sp}=67$ species and $N_{rxn}=263$ reactions. The forward reaction rates for each reaction are modeled as either; modified Arrhenius reactions ($k_{f}=AT^{b}exp(-E_{a}/R_{u}T)$, three-body reactions, or Troe falloff reactions. Each of these forward reaction rates are explained in more detail in [43]. The constants used by each reaction are included as supplementary material to this paper in a Cantera [44] yaml file format or can be found in CHEMKIN [45] formatted files in the original study [42]. $\vec{V}_{T}$ is the thermophoretic velocity associated with soot ($C(s)$) thermophoresis [46],
$\vec{V_{T}}=\frac{\mu}{2\rho T}\vec{\nabla}T.$ (4)
A six step semi-empirical soot model based on an acetylene precursor is adapted into the chemical mechanism that includes reactions for nucleation, agglomeration, surface growth and oxidation, which are given in full detail in [46].
[…]
$\mu(T)=\mu_{o}\left(\frac{T}{T_{o}}\right)^{3/2}\frac{T_{o}+S_{o}}{T+S_{o}}.$ (5)
The species diffusivity coefficients are assumed to be equal for all gaseous species. The thermal conductivity, $k$, and gas species diffusivity, $D$, are determined from $\mu$, assuming a unity Lewis number and Prandtl number of 0.707. The soot species diffusivity is assumed to be mostly from thermophoresis, $D_{Y_{C(s)}}\vec{\nabla}Y_{C(s)}=Y_{C(s)}\vec{V_{T}}$.
@@
The wall boundary conditions are simply treated as isothermal no-slip walls held at the ambient temperature of 300 K. The inlet and outlet boundary conditions are described using locally one-dimensional inviscid characteristics (LODI) [49]. The freestream velocity profile prescripted at the inlet is described using a $1/7^{th}$ power law for fully developed turbulent profile for pipe flow, $u_{in}(r)=V_{in}(1-2r/D_{pipe})^{1/7}$, where $V_{in}=60~{}G_{ox}/49\rho_{ox}$ is a velocity factor defined by the oxidizer mass flux ($G_{ox}$) to obtain the correct mass flow rate, the oxidizer density is $\rho_{ox}=1.283\frac{kg}{m^{3}}$, and the inlet is $D_{pipe}=0.0254m$.
@@
where $\dot{q}^{\prime\prime}_{c,g}(=k\vec{\nabla}T\cdot\hat{n}_{s})$ is the conductive heat flux from the gas phase normal to the fuel surface, $\dot{q}^{\prime\prime}_{c,f}$ is the conductive heat flux used in heating the fuel, and $\dot{q}^{\prime\prime}_{r,Net}(=\alpha_{s}\dot{q}^{\prime\prime}_{r,abs}-\epsilon_{s}\sigma T_{f}^{4}$) is the net radiation heat transfer at the fuel surface. $\dot{q}^{\prime\prime}_{r,abs}$ is the irradiant radiation heat flux onto the fuel surface, and $T_{f}$ is the fuel surface temperature. $\epsilon_{s}=0.9$ and $\alpha_{s}=0.9$ are the fuel surface emissivity and absorptivity. $\dot{m}^{\prime\prime}_{f}=\rho_{f}\dot{r}$ is the local fuel mass flux and is related to the local regression rate by the solid fuel density $\rho_{f}=1190\frac{kg}{m^{3}}$. The current simulations assume the fuel is isothermal and already at the vaporization temperature of PMMA ($T_{f}=653~{}K$) such that the is no heat loss due to heating the fuel ($\dot{q}_{c,f}=0$). This leads to the following mass flux and regression rate relations depending only on the instantaneous gas phase solution,
@@
In this isothermal fuel boundary model, the fuel temperature is held at $T_{f}=653K$ and the latent heat of sublimation at $l_{v}=840890J/kg$. The density, MMA species mass, and normal momentum boundary conditions have additional mass and momentum fluxes due to $\dot{m}^{\prime\prime}_{f}$.
@@
Where $\mathcal{S}_{B}$ includes the paired coordinates $x,y$ for the grid points on the boundary, $N$ is the number of simulation time steps, $t_{i}$ is the elapsed simulation time, and $\dot{r}(x,y,t)$ is the regression rate as computed during the simulation from Eqn. 7.
[…]
Figure 2: Relative error in $\tilde{\tilde{\dot{r}}}$ as a function of the grid step size compared to the finest mesh size case ($\Delta x$ = 0.05 $mm$)